# A Survey for Foundation Models in Autonomous Driving

### v3

@@ 3.1 Perception
SAM3DZhang et al. (2023a) applies SAM(Segment-anything model) to 3D object detection in autonomous driving. Lidar point clouds are projected to BEV(bird-eye-view) images, and it uses 32x32 mesh grids to generate point prompts to detect masks for foreground objects. It leverages the SAM model’s zero-shot transfer capability to generate segmentation masks and 2D boxes. Then it uses vertical attributes of those lidar points inside 2D boxes to generate 3D boxes. However, the Waymo Open Dataset evaluation shows the average-precision metrics are still far from existing state-of-the-art 3D object detection models. They observed that SAM trained foundation model can not handle those sparse and noisy points very well, and often results in false negatives for distant objects.
SAM is applied to domain adaptation for 3D segmentation tasks, leveraging the SAM model’s feature space which contains more semantic information and generalization capability. Peng et al. (2023) proposes SAM-guided feature alignment, learning unified representation of 3D point cloud features from different domains. It uses the SAM feature extractor to generate the camera image’s feature embedding and projects 3D point clouds into camera images to obtain SAM features. The training process optimizes the alignment loss so that 3D features from different domains have a unified representation in SAM’s feature space. This approach achieves state-of-the-art performance in 3D segmentation in multiple domain switching datasets, e.g. different cities, weathers, and lidar devices.
SAM and Grounding-DINOLiu et al. (2023b) are used to create a unified segmentation and tracking framework leveraging temporal consistency between video framesCheng et al. (2023). Grounding-DINO is an open-set object detector that takes input from text descriptions of objects and outputs the corresponding bounding boxes. Given the text prompts of object classes related to autonomous driving, it can detect objects in video frames and generate bounding boxes of vehicles and pedestrians. SAM model further takes these boxes as prompts and generates segmentation masks for detected objects. The resulting masks of objects are then passed to the downstream tracker, which compares the masks from continuous frames to determine if there are new objects.
@@ 3.2 Video Generation and World Model
GAIA-1Hu et al. (2023) is developed by Wayve to generate realistic driving videos. The world model uses camera images, text descriptions, and vehicle control signals as input tokens and predicts the next frame. The paper uses pre-trained DINOCaron et al. (2021) model’s embedding and cosine similarity loss to distill more semantic knowledge to image token embedding. They use the video diffusion modelHo et al. (2022) to decode high-fidelity driving scenes from the predicted image token. There are two separate tasks to train the diffusion model: image generation and video generation. The image generation task helps the decoder generate high-quality images, while the video generation task uses temporal attention to generate temporally consistent video frames. The generated video follows high-level real-world constraints and has realistic scene dynamics, such as the object’s location, interactions, traffic rules, and road structures. The video also shows diversity and creativity, which have realistic possible outcomes conditioned on different text descriptions and the ego vehicle’s action.
DriveDreamerWang et al. (2023b) also uses the world model and diffusion model to generate video for autonomous driving. In addition to images, text descriptions, and vehicle actions, the model also uses more structural traffic information as input, such as HDMap and object 3D boxes, so that the model can better understand higher-level structural constraints of traffic scenes. The model training has two stages: the first stage is video generation using the diffusion model conditioned on structured traffic information. It was built on a pre-trained Stable-Diffusion modelRombach et al. (2022) with parameters frozen. In the second stage, the model is trained with both future video prediction tasks and action prediction tasks to better learn future prediction and interactions between objects.
Zhang et al. (2023c) built a point cloud-based world model that achieves SOTA performance in point cloud forecasting tasks. They propose a VQVAE-likeOord et al. (2017) tokenizer to represent 3D point clouds as latent BEV tokens and use discrete diffusion to forecast future point clouds given the past BEV tokens and ego vehicle’s actions tokens.
@@ 4 Multi-modal Foundation Models
One of the most well-known multi-modal foundation models is CLIPRadford et al. (2021). The model is pre-trained using the contrastive pre-training method. The inputs are noisy images and text pairs, and the model is trained to predict if the given image and text are a correct pair. The model is trained to maximize the cosine similarity of embedding from the image encoder and text encoder. The CLIP model shows zero-shot transfer ability for other computer vision tasks, such as image classification, and predicting the correct text description of the class without supervised training.
Multi-modal foundation models, like LLaVALiu et al. (2023a), LISALai et al. (2023), and CogVLMWang et al. (2023a) can be used for the general-purpose visual AI agent, which demonstrates superior performance in vision tasks, such as object segmentation, detection, localization, and spatial reasoning. Video-LLaMAZhang et al. (2023b) can further perceive video and audio data, which may help autonomous vehicles better understand the world from temporal images and audio sequences.
Multi-modal foundation model is also used for robot learning, which leverages the robot’s action as a new modality to create more general-purpose agents that can perform real-world tasks. DeepMind proposed a vision-language-action modelBrohan et al. (2023) trained on text and images from the web and learned to output control commands to complete real-world object manipulation tasks.
@@ 4.1 Visual Understanding and Reasoning
To further improve the perception system, HiLM-DDing et al. (2023) utilizes multi-modal foundation models for ROLISP(Risk Object Localization and Intention and Suggestion Prediction). It uses natural language to identify risky objects from camera images and provide suggestions on the ego vehicle’s actions. To overcome the drawback of missing small objects, it proposes a pipeline with both high-resolution and low-resolution branches. The low-resolution reasoning branch is used to understand high-level information and identify risk objects from continuous video frames; The high-resolution perception branch enables further refinement of object detection and localization quality. Their model backbone uses the pre-trained visual encoder and LLM weights following BLIP2Li et al. (2023).
Talk2BEVDewangan et al. (2023) proposes an innovative bird’s-eye view (BEV) representation of the scene fusing both visual and semantic information. The pipeline first generates the BEV map from image and lidar data and uses general-purpose visual-language foundation models to add more detailed text descriptions of cropped images of objects. The JSON text representation of the BEV map is then passed to general-purpose LLM to perform Visual QA, which covers spatial and visual reasoning tasks. The result shows a good understanding of detailed instance attributes and also higher-level intent of objects, and the ability to provide free-formed advice on the ego vehicle’s actions.
LiDAR-LLMYang et al. (2023a) uses a novel approach that combines point cloud data with the advanced reasoning abilities of Large Language Models to interpret real-world 3D environments and achieves excellent performance in 3D captioning, grounding, and QA tasks. The model employs a unique three-stage training and a View-Aware Transformer(VAT) to align 3D data with text embedding, enhancing spatial comprehension. Their examples show the model can understand the traffic scenes and provide suggestions for autonomous driving planning tasks.
Atakishiyev et al. (2023) focus on the the explainability of vehicle’s actions using a visual QA approach. They collected driving videos in simulated environments from 5 different action categories(like going straight and turning left) and used manually labeled explanations of actions to train the model. The model was able to explain the driving decision based on road geometry and clearance of obstacles. They find it promising to apply state-of-the-art multi-modal foundation models to generate structured explanations of vehicle actions.
@@ 4.2 Unified Perception and Planning
Wen et al. (2023b) performed an early exploration of GPT-4VisionAchiam et al. (2023)’s application in perception and planning tasks, and evaluated its capabilities in several scenarios. It shows that GPT-4Vision can understand weather, traffic signs, and traffic lights and identify traffic participants in the scene. It can also provide more detailed semantic descriptions of these objects, such as vehicle rear lights, intents like U-turn, and detailed vehicle types(e.g. cement mixer truck, trailer, and SUV). It also shows the foundation model’s potential for understanding point cloud data, GPT-4V can identify vehicles from point cloud contours projected in BEV images. They also evaluated the model’s performance on planning tasks. Given the traffic scenario, GPT4-V is asked to describe its observation and decision on the vehicle’s action. The results show good interaction with other traffic participants and compliance with the traffic rules and common sense, e.g. following the car at a safety distance, yielding to cyclists at a crosswalk, remaining stopped until the light turns green. It can even handle some long-tail scenarios very well, such as the gated parking lot.
Instruction tuning is used to better adapt general-purpose multi-modal foundation models to autonomous driving tasks. DriveGPT4Xu et al. (2023) created an instruction-following dataset, where ChatGPT, YOLOV8Reis et al. (2023) and ground truth vehicle control signals from BBD-X datasetKim et al. (2018) are used to generate question and answers about common objects detections, spatial relations, traffic light signals, the ego vehicle’s actions. Following LLaVA, it used the pre-trained CLIPRadford et al. (2021) encoder and LLM weights and fine-tuned the model with their instruction-following dataset specifically designed for autonomous driving. They were able to build an end-to-end interpretable autonomous driving system, which is able to have a good understanding of the surrounding environment and make decisions on vehicle actions with jurisdictions and lower-level control commands.
@@ 4.3 Limitations and Future Directions
There are still limitations of multi-foundation models in autonomous driving. Wen et al. (2023b) shows the GPT-4V model still suffers from hallucination and generates unclear responses or false answers in several examples. The model also shows incompetence in utilizing multi-view cameras and lidar data for accurate 3D object detections and localization, because the pre-training dataset only contains 2D images from the web. More domain-specific fine-tuning or pre-training is required to train multi-modal foundation models to better understand point cloud data and sensor fusion to achieve comparable performance of the state-of-the-art perception system.
@@ 5 Conclusion and Future Directions
We also notice that the dataset is one of the biggest obstacles in the future development of foundation models in autonomous driving. The existing open-sourced datasetLi et al. (2024) for autonomous driving at the scale of 1000 hours, is far less than pre-training datasets used for state-of-the-art LLMs. The web dataset used for existing foundation models doesn’t leverage all modalities required by autonomous driving, such as lidar and surround cameras. The web data domain is also quite different from the real driving scenes.

### v4

@@ 3.1 Perception
SAM3D [39] applies SAM(Segment-anything model) to 3D object detection in autonomous driving. Lidar point clouds are projected to BEV(bird-eye-view) images, and it uses 32x32 mesh grids to generate point prompts to detect masks for foreground objects. It leverages the SAM model’s zero-shot transfer capability to generate segmentation masks and 2D boxes. Then it uses vertical attributes of those lidar points inside 2D boxes to generate 3D boxes. However, the Waymo Open Dataset evaluation shows the average-precision metrics are still far from existing state-of-the-art 3D object detection models. They observed that SAM trained foundation model can not handle those sparse and noisy points very well, and often results in false negatives for distant objects.
SAM is applied to domain adaptation for 3D segmentation tasks, leveraging the SAM model’s feature space which contains more semantic information and generalization capability. [40] proposes SAM-guided feature alignment, learning unified representation of 3D point cloud features from different domains. It uses the SAM feature extractor to generate the camera image’s feature embedding and projects 3D point clouds into camera images to obtain SAM features. The training process optimizes the alignment loss so that 3D features from different domains have a unified representation in SAM’s feature space. This approach achieves state-of-the-art performance in 3D segmentation in multiple domain switching datasets, e.g. different cities, weathers, and lidar devices.
SAM and Grounding-DINO[41] are used to create a unified segmentation and tracking framework leveraging temporal consistency between video frames[42]. Grounding-DINO is an open-set object detector that takes input from text descriptions of objects and outputs the corresponding bounding boxes. Given the text prompts of object classes related to autonomous driving, it can detect objects in video frames and generate bounding boxes of vehicles and pedestrians. SAM model further takes these boxes as prompts and generates segmentation masks for detected objects. The resulting masks of objects are then passed to the downstream tracker, which compares the masks from continuous frames to determine if there are new objects.
@@ 3.2 Video Generation and World Model
GAIA-1[43] is developed by Wayve to generate realistic driving videos. The world model uses camera images, text descriptions, and vehicle control signals as input tokens and predicts the next frame. The paper uses pre-trained DINO[28] model’s embedding and cosine similarity loss to distill more semantic knowledge to image token embedding. They use the video diffusion model[44] to decode high-fidelity driving scenes from the predicted image token. There are two separate tasks to train the diffusion model: image generation and video generation. The image generation task helps the decoder generate high-quality images, while the video generation task uses temporal attention to generate temporally consistent video frames. The generated video follows high-level real-world constraints and has realistic scene dynamics, such as the object’s location, interactions, traffic rules, and road structures. The video also shows diversity and creativity, which have realistic possible outcomes conditioned on different text descriptions and the ego vehicle’s action.
DriveDreamer [45] also uses the world model and diffusion model to generate video for autonomous driving. In addition to images, text descriptions, and vehicle actions, the model also uses more structural traffic information as input, such as HDMap and object 3D boxes, so that the model can better understand higher-level structural constraints of traffic scenes. The model training has two stages: the first stage is video generation using the diffusion model conditioned on structured traffic information. It was built on a pre-trained Stable-Diffusion model[33] with parameters frozen. In the second stage, the model is trained with both future video prediction tasks and action prediction tasks to better learn future prediction and interactions between objects.
[46] built a point cloud-based world model that achieves SOTA performance in point cloud forecasting tasks. They propose a VQVAE-like [47] tokenizer to represent 3D point clouds as latent BEV tokens and use discrete diffusion to forecast future point clouds given the past BEV tokens and ego vehicle’s actions tokens.
@@ 4 Multi-modal Foundation Models
One of the most well-known multi-modal foundation models is CLIP[48]. The model is pre-trained using the contrastive pre-training method. The inputs are noisy images and text pairs, and the model is trained to predict if the given image and text are a correct pair. The model is trained to maximize the cosine similarity of embedding from the image encoder and text encoder. The CLIP model shows zero-shot transfer ability for other computer vision tasks, such as image classification, and predicting the correct text description of the class without supervised training.
Multi-modal foundation models, like LLaVA[49], LISA[50], and CogVLM[51] can be used for the general-purpose visual AI agent, which demonstrates superior performance in vision tasks, such as object segmentation, detection, localization, and spatial reasoning. Video-LLaMA[52] can further perceive video and audio data, which may help autonomous vehicles better understand the world from temporal images and audio sequences.
Multi-modal foundation model is also used for robot learning, which leverages the robot’s action as a new modality to create more general-purpose agents that can perform real-world tasks. DeepMind proposed a vision-language-action model[53] trained on text and images from the web and learned to output control commands to complete real-world object manipulation tasks.
@@ 4.1 Visual Understanding and Reasoning
To further improve the perception system, HiLM-D[54] utilizes multi-modal foundation models for ROLISP(Risk Object Localization and Intention and Suggestion Prediction). It uses natural language to identify risky objects from camera images and provide suggestions on the ego vehicle’s actions. To overcome the drawback of missing small objects, it proposes a pipeline with both high-resolution and low-resolution branches. The low-resolution reasoning branch is used to understand high-level information and identify risk objects from continuous video frames; The high-resolution perception branch enables further refinement of object detection and localization quality. Their model backbone uses the pre-trained visual encoder and LLM weights following BLIP2[55].
Talk2BEV[56] proposes an innovative bird’s-eye view (BEV) representation of the scene fusing both visual and semantic information. The pipeline first generates the BEV map from image and lidar data and uses general-purpose visual-language foundation models to add more detailed text descriptions of cropped images of objects. The JSON text representation of the BEV map is then passed to general-purpose LLM to perform Visual QA, which covers spatial and visual reasoning tasks. The result shows a good understanding of detailed instance attributes and also higher-level intent of objects, and the ability to provide free-formed advice on the ego vehicle’s actions.
LiDAR-LLM[57] uses a novel approach that combines point cloud data with the advanced reasoning abilities of Large Language Models to interpret real-world 3D environments and achieves excellent performance in 3D captioning, grounding, and QA tasks. The model employs a unique three-stage training and a View-Aware Transformer(VAT) to align 3D data with text embedding, enhancing spatial comprehension. Their examples show the model can understand the traffic scenes and provide suggestions for autonomous driving planning tasks.
[58] focus on the the explainability of vehicle’s actions using a visual QA approach. They collected driving videos in simulated environments from 5 different action categories(like going straight and turning left) and used manually labeled explanations of actions to train the model. The model was able to explain the driving decision based on road geometry and clearance of obstacles. They find it promising to apply state-of-the-art multi-modal foundation models to generate structured explanations of vehicle actions.
@@ 4.2 Unified Perception and Planning
[59] performed an early exploration of GPT-4Vision[5]’s application in perception and planning tasks, and evaluated its capabilities in several scenarios. It shows that GPT-4Vision can understand weather, traffic signs, and traffic lights and identify traffic participants in the scene. It can also provide more detailed semantic descriptions of these objects, such as vehicle rear lights, intents like U-turn, and detailed vehicle types(e.g. cement mixer truck, trailer, and SUV). It also shows the foundation model’s potential for understanding point cloud data, GPT-4V can identify vehicles from point cloud contours projected in BEV images. They also evaluated the model’s performance on planning tasks. Given the traffic scenario, GPT4-V is asked to describe its observation and decision on the vehicle’s action. The results show good interaction with other traffic participants and compliance with the traffic rules and common sense, e.g. following the car at a safety distance, yielding to cyclists at a crosswalk, remaining stopped until the light turns green. It can even handle some long-tail scenarios very well, such as the gated parking lot.
Instruction tuning is used to better adapt general-purpose multi-modal foundation models to autonomous driving tasks. DriveGPT4 [60] created an instruction-following dataset, where ChatGPT, YOLOV8 [61] and ground truth vehicle control signals from BBD-X dataset [62] are used to generate question and answers about common objects detections, spatial relations, traffic light signals, the ego vehicle’s actions. Following LLaVA, it used the pre-trained CLIP[48] encoder and LLM weights and fine-tuned the model with their instruction-following dataset specifically designed for autonomous driving. They were able to build an end-to-end interpretable autonomous driving system, which is able to have a good understanding of the surrounding environment and make decisions on vehicle actions with jurisdictions and lower-level control commands.
@@ 4.3 Limitations and Future Directions
There are still limitations of multi-foundation models in autonomous driving. [59] shows the GPT-4V model still suffers from hallucination and generates unclear responses or false answers in several examples. The model also shows incompetence in utilizing multi-view cameras and lidar data for accurate 3D object detections and localization, because the pre-training dataset only contains 2D images from the web. More domain-specific fine-tuning or pre-training is required to train multi-modal foundation models to better understand point cloud data and sensor fusion to achieve comparable performance of the state-of-the-art perception system.
@@ 5 Conclusion and Future Directions
We also notice that the dataset is one of the biggest obstacles in the future development of foundation models in autonomous driving. The existing open-sourced dataset[63] for autonomous driving at the scale of 1000 hours, is far less than pre-training datasets used for state-of-the-art LLMs. The web dataset used for existing foundation models doesn’t leverage all modalities required by autonomous driving, such as lidar and surround cameras. The web data domain is also quite different from the real driving scenes.
@@
Conflict of Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Availability of Data
The data that support the findings of this study are available from the corresponding author upon request.
Ethical Considerations
This study is a literature review and did not involve human participants or animal subjects. Therefore, formal ethics approval was not required.
Funding Information
This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Author Contributions
Y.Shen conceptualized the study. H. Gao, Z.Wang, Y. Li, and K. Long performed the literature review and analysis. H. Gao and Y. Shen contributed to the interpretation of the results. H. Gao and Y. Shen drafted the manuscript. M. Yang and Y. Shen contributed to the revision of the manuscript and approved the final version.